OpenStudy (anonymous):

please check my work (first one to check my work gets a medal) use substitution to solve the following system of equations d+e-f e=f+d+5 f=2e-4

OpenStudy (anonymous):

i got what got: e:-1 f:-6 d:-10

OpenStudy (anonymous):

Is that first equation written correctly ? Or is there supposed to be an equal sign in there ?

OpenStudy (anonymous):

my bad d+e-f=-5

OpenStudy (anonymous):

f = 2e - 4 sub this into the 1st equation d + e - f = - 5 d + e - (2e - 4) = - 5 d + e - 2e + 4 = - 5 d - e = -9 d = e - 9 now sub e - 9 in for d and 2e - 4 in for f in the 2nd equation e = f + d + 5 e = (2e - 4) +( e - 9) + 5 (combine like terms) e = 3e - 13 + 5 -2e = - 8 e = -8/-2 e = 4 now sub 4 in for e in the 3rd equation f = 2e - 4 f = 2(4) - 4 f = 8 - 4 f = 4 now sub 4 in for e and 4 in for f in the 1st equation d + e - f = - 5 d + 4 - 4 = - 5 d = - 5 check... e = f + d + 5 4 = 4 - 5 + 5 4 = 4 (correct) or d + e - f = - 5 -5 + 4 - 4 = - 5 -5 = -5 (correct) or f = 2e - 4 4 = 2(4) - 4 4 = 8 - 4 4 = 4 (correct) I got : d = -5, e = 4, and f = 4

OpenStudy (anonymous):

since your the only one that replied i'll give you the medal thanks for the answer too

OpenStudy (anonymous):

your quite welcome.....do you see how I did it ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

good, because I am not very good at explaining...lol :)

OpenStudy (anonymous):

yeah you are you just gotta believe in yourself :)

OpenStudy (anonymous):

sweet words spoken by the " wild " one.....thank you :)

OpenStudy (anonymous):

any time

OpenStudy (anonymous):

your the one with sweet words